OpenStudy (anonymous):

order from least to greatest. and please explane how to do this 4^2 x (3 + 5) ; 4 x √400 ; -4 x (4 - 5^2) x -2

OpenStudy (lgbasallote):

1) 4^2 x (3+5) = 16 x 8 = 128 2) 4 x \(\sqrt{400}\) = 4 x 20 = 80 3) -4 x (4 - 5^2) x -2 = -4 x(4 - 25) x -2= -4 x (-21) x -2 = 84 x -2 = -168 so now can you arrange them from least to greatest?

OpenStudy (anonymous):

yep..how do u find the √?

OpenStudy (lgbasallote):

\[\Large \sqrt{400} \implies \sqrt{4 \times 100} \implies \sqrt 4 \times \sqrt{100} \implies 2 \times 10 \implies 20\]

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

no.. how did the 4 and the 100 go to a 2 and a 10?

OpenStudy (lgbasallote):

square root of 4 is 2...square root of 100 is 10

OpenStudy (anonymous):

ok so then the 2 x 10 would go to a 20 x 4?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

so were did the last 4 come from again?

OpenStudy (lgbasallote):

4 times sqrt(400) sqrt 400 is 20 4 x 20 = 80

OpenStudy (anonymous):

kk thx

OpenStudy (anonymous):

this might help u how to get10 and 2 u root them both to get ur answer